Compute the expected age u of a Bristish nurse in 1851. Assume that the table below shows the age distribution of nurses in Great Britain in 1851. Round your answer to nearest hundredth.
Answer #1
Age range (yr)

20–29

30–39

40–49

50–59

60–69

70–79

80+

Midpoint (x)

24.5

34.5

44.5

54.5

64.5

75.5

84.5

Percent of nurses

5.6%

9.8%

19.4%

29.2%

25.1%

9.1%

1.8%

The Expected age is given by r*p(x)

= 24.5*5.6/100 + 34.5*9.8/100 + 44.5*19.4/100 + 54.5*29.2/100 + 64.5*25.1/100 + 75.5*9.1/100 + 84.5 * 1.8/100

= 1.372 + 3.381 + 8.633 + 15.914 + 16.1895 + 6.8705 + 1.521 = 53.881 approx53.88